|-25| - (-32)
-3.35 - 6.21

Found 2 solutions by checkley77, scottfin:
Answer by checkley77(12844)

|-25| - (-32)
-25+32=-7 ans.
+25+32=57 ans.
-3.35 - 6.21=-9.56 ans.

Answer by scottfin(2)

|-25|-(-32)-3.35-6.21
25-(-32)-3.35-6.25
25+32-3.35-6.21
57-3.35-6.21
53.65-6.21
47.44
The absolute value of -25 is 25. A negative times a negative is a positive, as in
-(-32)=32. Put the rest in a line, as in line 3, and finish solving. If one solves the way I do, you will end up with a upside down triangle by solving left to right, one operation at a time. On here, it looks like half a triangle.